# Wake Vision: A Large-scale, Diverse Dataset and Benchmark Suite for TinyML Person Detection

### v1

@@ Abstract
Machine learning applications on extremely low-power devices, commonly referred to as tiny machine learning (TinyML), promises a smarter and more connected world. However, the advancement of current TinyML research is hindered by the limited size and quality of pertinent datasets. To address this challenge, we introduce Wake Vision, a large-scale, diverse dataset tailored for person detection—the canonical task for TinyML visual sensing. Wake Vision comprises over 6 million images, which is a hundredfold increase compared to the previous standard, and has undergone thorough quality filtering. Using Wake Vision for training results in a 2.41% increase in accuracy compared to the established benchmark. Alongside the dataset, we provide a collection of five detailed benchmark sets that assess model performance on specific segments of the test data, such as varying lighting conditions, distances from the camera, and demographic characteristics of subjects. These novel fine-grained benchmarks facilitate the evaluation of model quality in challenging real-world scenarios that are often ignored when focusing solely on overall accuracy. Through an evaluation of a MobileNetV2 TinyML model on the benchmarks, we show that the input resolution plays a more crucial role than the model width in detecting distant subjects and that the impact of quantization on model robustness is minimal, thanks to the dataset quality. These findings underscore the importance of a detailed evaluation to identify essential factors for model development. The dataset, benchmark suite, code, and models are publicly available under the CC-BY 4.0 license, enabling their use for commercial use cases.
Keywords:
Dataset Person Detection TinyML Benchmarks
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
To reduce the energy consumption and cost of deploying Machine Learning (ML) deployments, researchers have pioneered a new field of ultra-low-power machine learning (TinyML) [3, 35, 1, 8]. TinyML aims to co-locate ML models with the sensors by deploying the model onto always-on microcontrollers (MCUs) and TinyML accelerators, thereby reducing the energy cost of transmitting data or constantly running larger processors. Deploying models on these tiny devices imposes severe constraints, typically confining the model’s size to a few hundred kilobytes, nearly four orders of magnitude less memory than mobile phones [2].
@@
The field of TinyML requires large-scale datasets to facilitate high-quality research. However, the extreme constraints imposed by the hardware necessitate that TinyML models be compact and efficient, rendering them unsuitable for traditional, complex ML tasks. As a result, conventional ML datasets like ImageNet [9], designed for more resource-intensive models, are often ill-suited for TinyML research endeavors. While existing TinyML datasets [7, 34] have been valuable resources, they tend to be limited in scale, making it challenging to train production-grade TinyML models solely on these datasets.
@@
To enable and mature the field of TinyML research, we present the Wake Vision dataset, a large dataset for person detection, the quintessential vision use case for TinyML [1]. Person detection is a binary image classification task where a model detects whether a person is present. While simple, this task has many important use cases, such as occupancy detection [29], smart HVAC/lighting [36], or acting as an always-on ‘wake model’ in a larger ML system [14].
Wake Vision has over 6M images, ~100x more than the prior state-of-the-art person detection dataset, Visual Wake Words (VWW) [7]. Wake Vision is derived from the Open Images v7 dataset [21, 19] and is permissively licensed. Wake Vision demonstrates a 2.41% accuracy improvement over VWW.
Due to the limited capacity of TinyML models, data quality is critical [26], however techniques, such as knowledge distillation [15], can leverage large, noisy datasets more effectively. For this reason, we choose to release two training sets: one that prioritizes dataset size, Wake Vision (Large), and one that prioritizes data quality, Wake Vision (Quality). We find that Wake Vision (Quality) out performs Wake Vision (Large), but we release both training sets to enable research into the tradeoffs between quantity and quality of data for TinyML.
@@
Person detection systems are deployed, and expected to work, in challenging settings, such as low lighting. However, because these challenging settings are uncommon in images sourced from the internet, high-level metrics, like test set accuracy, may overestimate a model’s real world performance. To address this issue we introduce a suite of five fine-grain benchmarks, which contain only images of a particular demographic or scenario. Our benchmarks cover distance, lighting, depictions, and perceived gender and age. Finally, we use our benchmark suite to conduct a case study on the performance effects of typical TinyML compression. This analysis demonstrates the usefulness of our benchmark and can inform future TinyML model design.
In summary, we present the following novel contributions:
@@
  • A new person detection dataset, called Wake Vision, which is an easily accessible, permissively licensed dataset that is ~100x larger than the prior standard.
@@
  • A suite of five fine-grain benchmark sets that evaluate a model’s fairness and robustness in challenging settings.
  • A case study on the effect of TinyML scaling the highlights the importance of fine grain analysis.
@@ 2 Related Work
The current de facto standard dataset for person detection is the VWW dataset [7, 8, 23, 2, 1]. To our knowledge, this is the only current open-source dataset explicitly meant for person detection. Unfortunately, the VWW dataset is small and cannot be directly downloaded. Instead, it must be regenerated from MS-COCO [24], thus limiting its accessibility and usefulness.
Apart from dedicated person detection datasets, some general image classification datasets contain a person label, which makes it possible to create person detection models based on them. Examples include the Cifar-100 dataset [20], and the PASCAL Visual Object Classes dataset [11]. However, using general image classification datasets can lead to poor perceived performance of TinyML models, as they do not adequately represent the open “no-person” class [7].
Tab. 1 shows that our Wake Vision dataset contains up to almost two orders of magnitude more images than any other public and permissively licensed dataset. Furthermore, Wake Vision is the only person detection dataset to provide a fine-grained benchmark suite and be officially distributed to popular dataset services such as TensorFlow Datasets (TFDS) and Hugging Face Datasets.
Furthermore, Wake Vision integrates fine-grained benchmarks, which are a novel addition to existing vision data sets for TinyML. These fine-grained benchmark sets, which include distance, lighting, depictions, and perceived gender and age variations, allow for a deeper investigation into the effects of input and model compression on specific demographic or scenario subsets. Several research papers report little performance degradation from model compression on top-line metrics such as top-1 test set accuracy [6, 13, 10, 18]. However, recent research suggests that this focus on top-line metrics can hide poor performance on critical subsets of a dataset, e.g., subsets relating to sensitive information and fairness [16, 17].
@@
Total # of Person Images Fine-Grain Suitable for
@@
Wake Vision (Quality) 1,423,690 675,411 9,106 27,328
Wake Vision (Large) 6,477,906 3,238,953 - -
@@ 3 Wake Vision Dataset
We introduce the Wake Vision dataset, the largest dataset for TinyML person detection. The dataset’s size, quality, and detailed metadata enable new avenues of TinyML research. Wake Vision is two orders of magnitude larger than prior comparable datasets and can be adapted based on the target use case. Fig. 1 illustrates examples of images in the dataset. The dataset is permissively licensed for research and commercial use. Images are sourced from Flickr [12] images listed as having a CC-BY 2.0 license Labels are derived from Open Images and licensed under the CC-BY 4.0 license. In this section, we discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the dataset and directly comparing it to the Visual Wake Dataset.
Older Person
Near Person
Bright Image
Female Person
Depicted Person
Young Person
@@ 3.1 Label Generation
While a sufficiently large person detection dataset is indispensable for TinyML research, it can be challenging to obtain. Manually labelling millions of images would be prohibitively expensive for a nascent research field like TinyML.
Therefore, we leverage existing large-scale data efforts to to generate Wake Vision automatically. The base label in Wake Vision is a binary person/non-person label. We derive these labels from Open Images, which contain both image-level and bounding box labels. Image-level labels describe which objects are present in an image but not localized to a specific point in the image.
An image can contain several image-level labels, where as bounding box labels provide a set of four coordinates forming a box around the object. The bounding box label classes of Open Images are hierarchically structured so that one class can be a subcategory or a part of another class, e.g., a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.”
We observe that the bounding box labels are generally more accurate and provide more information for data filtering but are less numerous. The Open Images training set has over 3 million images with image-level person labels, but only ~ 700,000 images with a person bounding box. The Open Images validation and test sets are fully labeled with image-level and bounding box labels. Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels), as different use cases may have different requirements for the training data (e.g., active learning).
We provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality). One larger set, labeled via Open Images image-level labels, and one smaller set, labeled via Open Images bounding box labels (Tab. 1). In our experiments, the Wake Vision (Quality) training set produce more accurate models than the Wake Vision (Large) training set, strongly suggesting that label quality is more important than quantity in our setting.
3.1.1 Large Scale Training Set from Image-Level Labels:
Image-level labels in Open Images contain little additional information in addition to the labels. The only additionally relevant information is a confidence property, which represents how certain it is that a label is correct. The confidence ranges from 0-1 for labels generated by machines, and is either 0 or 1 for labels negatively or positively verified by humans. We use this property to exclude images from Wake Vision if they contain only low confidence person labels.
3.1.2 High Quality Training Set From Bounding Box Labels:
@@
We allow users to adapt the Wake Vision dataset to meet the needs of specific usecases by adding several configurable options in the open-source dataset creation script. A comprehensive list of Open Images labels related to Wake Vision can be found in Appendix 0.I.
@@
3.1.3 Training Set Comparison
To compare the performance of the two training dataset sets, we train a MobileNetV2-0.25 [31] model on each dataset for an equal number of steps and compare its performance on the common test set (Tab. 2). Given the amount of label errors in the Wake Vision test set shown in Tab. 4, a perfect person detection model would achieve an accuracy of ~93%. The filtered training set generated from the bounding box labels outperforms the larger training set generated from the image-level labels by 2.65% test accuracy. This indicates that label quality is more important than quantity in this setting. Interestingly, although the gap between the two is reduced to just 0.36% when training exclusively on soft labels from a teacher model, the bonding box training set is still superior. This indicates that the difference in quality extends beyond just the labels and that our data filtering leads to an improvement on it’s own.
Table 2: Accuracy on Wake Vision Test set of a MobileNetV2-0.25 model trained on image-level and bounding box labels, with and without distillation. A MobileNetV2-1.0 model trained on the bonding box set is used as the teacher in distillation.
Training set Label Source Dataset Size Accuracy Distilled Accuracy
Wake Vision (Large) Image-Level 6,477,906 86.00% 88.96%
Wake Vision (Quality) Bounding Box 1,350,822 88.65% 89.32%
Despite the worse performance in this setting, the larger image-level training set may be useful for training higher capacity models for longer or when doing active learning [30] or distillation [5]. As our distillation results indicate, further data-centric filtering techniques may be needed to fully leverage the image-level training set. We experimented with tripling the model size and doubling the training time, but the difference in accuracy remained relatively consistent.
@@ 3.1.3 Training Set Comparison
3.2 Error Correction
Labeling errors are a challenge in computer vision [28, 4] that limit progress and obscures true performance. This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs. Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of its labels, aligning with best practices in the field of computer vision.
@@ 3.2 Error Correction
Given the substantial scale of Wake Vision, manually correcting label errors across the entire dataset becomes an arduous undertaking. To address this challenge, we employ the Confident Learning technique [27] to intelligently identify potential label errors. We selected Confident Learning as prior work has demonstrated its capability to find label errors in large datasets [28]. Once these suspected mislabeled instances are flagged, we methodically inspect and correct them through a manual verification process.
Dataset Total Suggested Corrected Suggestion
Split Size Errors Errors Accuracy
Validation 16,964 632 81 12.82%
Test 51,282 1672 267 15.97%
As shown in Tab. 3, the confident learning process identified a large amount of possible label errors in Wake Vision’s Validation and Test sets. However, only between 12 and 16% of these possible label errors were legitimate errors. We correct these identified label errors in the Wake Vision validation and test sets.
Given this low acceptance rate of label issues identified by Confident Learning, we conclude that we cannot automate label cleaning to a point where no human in the loop is needed. Based on that we also conclude that this strategy too human-intensive to be applied to the much larger training sets.
We report our estimated label error rate after label corrections versus VWW in Tab. 4. We find that after our corrections, the VWW dataset has a top-level error rate of 7.8%, which is comparable to Wake Vision’s top-level error rate of 6.8%. Notably, about a third of the Wake Vision label errors stem from depictions missing the depiction flag. As we by default do not consider depictions as persons, missing this flag results in a label error. Not considering the depiction label errors, Wake Vision has closer to half the number of errors of VWW.
VWW does not, to the authors knowledge, include depictions. Thus no depiction errors are present for the VWW dataset. In real world settings, depictions are everywhere, therefore we expect that Wake Vision will prove to be a more realistic dataset. ML systems based on Wake Vision will therefore likely be able to handle real world scenarios better than equivalent systems based on VWW.
We are in the process of crowd-sourcing label corrections for the validation and test sets to eliminate the remaining errors. However, despite these efforts, Wake Vision, in its current state, already offers a significant improvement over the VWW dataset, as detailed in Section 3.4. Although further enhancements to label quality are underway, Wake Vision presents a valuable and directly improved resource compared to its predecessor VWW dataset.
Table 4: Percentage (and count) of label errors in a 500 sample subset of the VWW and Wake Vision test sets. Percent of label errors caused by misclassified depictions are reported.
Dataset Label Error Rate % Errors From Depictions
Visual Wake Words 7.8% (39) 0% (0)
Wake Vision 6.8% (34) 38.2% (13)

### v2

# Wake Vision: A Large-scale, High-Quality Dataset and Benchmark Suite for TinyML Applications

@@ Abstract
Tiny machine learning (TinyML), which enables machine learning applications on extremely low-power devices, suffers from limited size and quality of relevant datasets. To address this issue, we introduce Wake Vision, a large-scale, diverse dataset tailored for person detection—the canonical task for TinyML visual sensing. Wake Vision comprises over 6 million images, representing a hundredfold increase compared to the previous standard, and has undergone thorough quality filtering. We provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality), a smaller set with higher-quality labels. Our results demonstrate that using the Wake Vision (Quality) training set produces more accurate models than the Wake Vision (Large) training set, strongly suggesting that label quality is more important than quantity in our setting. We find use for the large training set for pre-training and knowledge distillation. To minimize label errors that can obscure true model performance, we manually label the validation and test sets, improving the test set error rate from 7.8% in the prior standard to only 2.2%. In addition to the dataset, we provide a collection of five detailed benchmark sets to facilitate the evaluation of model quality in challenging real-world scenarios that are often ignored when focusing solely on overall accuracy. These novel fine-grained benchmarks assess model performance on specific segments of the test data, such as varying lighting conditions, distances from the camera, and demographic characteristics of subjects. These benchmarks enable researchers and practitioners to better understand the strengths and limitations of their models in real-world scenarios. Our results demonstrate that using Wake Vision for training results in a 2.49% increase in accuracy compared to the established dataset. We also show the importance of dataset quality for low-capacity models and the value of dataset size for high-capacity models. The dataset, benchmark suite, code, and models are publicly available under the CC-BY 4.0 license at https://wakevision.ai/.
**footnotetext: These authors contributed equally to this work.11footnotetext: Work done while a visiting researcher at Harvard University.
@@ 1 Introduction
To reduce the energy consumption and cost of deploying Machine Learning (ML) deployments, researchers have pioneered ultra-low-power machine learning (TinyML) [3, 30, 1, 6]. TinyML co-locates ML models with the sensors by deploying the model onto always-on microcontrollers (MCUs) and TinyML accelerators, thereby reducing the energy cost of transmitting data or constantly running larger processors. Deploying models on these tiny devices imposes severe constraints, confining the model’s size to $\sim$100 kilobytes, nearly four orders of magnitude less memory than mobile phones [2].
@@
TinyML requires large-scale datasets to facilitate high-quality research. However, the extreme constraints imposed by the hardware necessitate that TinyML models be compact and efficient, rendering them unsuitable for traditional, complex ML tasks. As a result, conventional ML datasets like ImageNet [7], designed for more resource-intensive models, are often ill-suited for TinyML research endeavors. While existing TinyML datasets [5, 29] have been valuable, they tend to be limited in scale, making it challenging to train production-grade TinyML models on these datasets.
@@
To enable and mature the field of TinyML research, we present the Wake Vision dataset, a large and high-quality dataset for person detection, the quintessential vision use case for TinyML [1]. Person detection is a binary image classification task where a model detects whether a person is present. While simple, this task has many important use cases, such as occupancy detection [25], smart HVAC/lighting [31], or acting as an always-on ‘wake model’ in a larger ML system [11]. Wake Vision has around  6M images, ~100x more than the prior state-of-the-art person detection dataset, Visual Wake Words (VWW) [5]. Wake Vision is derived from the Open Images v7 dataset [16, 14] and is permissively licensed. Wake Vision demonstrates a 2.21% accuracy improvement over VWW.
We leverage the Wake Vision dataset for an insightful study on the trade-offs between data quality and quantity in TinyML. We connect the dataset’s unique characteristics with the limitations of TinyML models. Prior work suggests that large, overparameterized models, which have more learnable parameters than data samples, can identify and adapt to errors in their training sets [32, 9]. However, TinyML models are rarely overparameterized, leading us to hypothesize that the trade-off between data quality and quantity may be skewed towards quality for these much smaller ML models.
To explore this hypothesis and showcase the value of the Wake Vision dataset in the TinyML domain, we release two training sets as part of the Wake Vision visual corpus: Wake Vision (Large), which prioritizes dataset size, and Wake Vision (Quality), which prioritizes data quality. Using a naive training script, we find that Wake Vision (Quality) outperforms Wake Vision (Large). However, our best-performing Wake Vision model is created by using the two sets in unison, employing Wake Vision (Large) for pre-training and Wake Vision (Quality) for fine-tuning. This approach demonstrates the potential of the Wake Vision dataset to advance TinyML research by leveraging both data quantity and quality. In addition to presenting these findings, we conduct several experiments designed to investigate the trade-offs between dataset quality and quantity across a range of models. To facilitate further research into these trade-offs for TinyML, we release both training sets.
@@
TinyML person detection systems are deployed and expected to function effectively in challenging settings, such as low lighting conditions. However, these challenging settings are often underrepresented in images sourced from the internet, which can lead to high-level metrics, like overall test set accuracy, overestimating a model’s real-world performance. To address this issue, we introduce a suite of five fine-grained benchmarks, each containing images of a specific demographic or scenario.
The benchmarks are designed to provide a more comprehensive evaluation of a person detection model’s performance across a range of real-world conditions. Our benchmarks cover several factors, including distance, which evaluates the model’s performance in detecting people at various distances from the camera; lighting, which assesses the model’s accuracy in detecting individuals in both well-lit and poorly-lit environments; depictions, which measures the model’s ability to handle different representations of people, such as drawings or sculptures; perceived gender, which investigates potential biases in the model’s performance when detecting people of different perceived genders; and age, which examines the model’s effectiveness in detecting individuals across various age groups.
By evaluating models using these fine-grained benchmarks contained within Wake Vision, we can identify potential biases, limitations, and areas for improvement that may not be apparent when relying solely on high-level metrics. This approach ensures that the person detection systems we develop are not only accurate on standard test sets but also robust and reliable across a diverse range of real-world settings and demographics. In summary, we present the following novel contributions:
@@
  • A new person detection dataset, called Wake Vision, which is an easily accessible, permissively licensed dataset that is two orders of magnitude larger than the prior standard.
  • An improved, high-quality test set that reduces the estimated label error rate to only 2.2% from the prior standard of 7.8%
@@
  • A suite of five fine-grained benchmark sets that evaluate a model’s fairness and robustness in challenging settings.
@@ 2 Related Work
The current de facto standard dataset for person detection is the VWW dataset [5, 6, 18, 2, 1]. To our knowledge, this is the only current open-source dataset designed explicitly for person detection. Unfortunately, the VWW dataset is small and cannot be directly downloaded. Instead, it must be regenerated from MS-COCO [19], thus limiting its accessibility and usefulness. Apart from dedicated person detection datasets, some general image classification datasets contain a person label, which makes it possible to create person detection models based on them. Examples include the Cifar-100 dataset [15], and the PASCAL Visual Object Classes dataset [8].
However, using general image classification datasets can lead to poor perceived performance of TinyML models, as they do not adequately represent the open “no-person” class [5]. Table 1 shows that our Wake Vision dataset contains up to almost two orders of magnitude more images than any other public and permissively licensed dataset. Furthermore, Wake Vision is the only person detection dataset to provide a fine-grained benchmark suite and be officially distributed to popular dataset services such as TensorFlow Datasets (TFDS) and Hugging Face Datasets. However, recent research suggests that this focus on top-line metrics can hide poor performance on critical subsets of a dataset, such as subsets relating to sensitive information and fairness [12, 13].
@@
Total # of Person Images Fine-Grained Suitable for
@@
Wake Vision (Quality) 1,322,574 624,115 9,2091 27,881
Wake Vision (Large) 5,760,428 2,880,214 - -
@@ 3 Wake Vision Dataset
We introduce the Wake Vision dataset, the largest dataset for TinyML person detection. The dataset’s size, quality, and detailed metadata enable new avenues of TinyML research. Wake Vision is two orders of magnitude larger than prior comparable datasets and can be adapted based on the target use case. We also demonstrate how its quality is significantly improved compared to existing datsets. Fig. 1 illustrates examples of images in the dataset. We discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the dataset and comparing it to the Visual Wake Words Dataset.
(a) Older Person
(b) Near Person
(c) Bright Image
(d) Female Person
(e) Depicted Person
(f) Young Person
@@ 3.1 Label Generation
While a large person detection dataset is indispensable for TinyML research, it can be challenging to obtain. Manually labelling millions of images would be prohibitively expensive for a nascent field like TinyML. Therefore, we use existing large-scale data efforts to generate Wake Vision automatically.
The base label in Wake Vision is a binary person/non-person label. We derive these labels from Open Images, which contain both image-level and bounding box labels. Image-level labels describe which objects are present in an image but not localized to a specific point in the image. Bounding box labels provide a set of four coordinates forming a box around the object. The bounding box label classes are hierarchically structured so that one class can be a subcategory or a part of another class. For example a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.”
We observe that the bounding box labels are generally more accurate and provide more information for data filtering but are less numerous. The Open Images training set has ~9 million images with image-level person labels, but only ~1.7 million images with a person bounding box. The Open Images validation and test sets are fully labeled with image-level and bounding box labels. Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels). Because of this trade-off, we provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality). One larger set, labeled via Open Images image-level labels, and one smaller set, labeled via Open Images bounding box labels (Table 1).
Large Training Set from Image-Level Labels
Image-level labels in Open Images contain little additional information in addition to the labels. The only additionally relevant information is a confidence property, which represents how certain it is that a label is correct. The confidence ranges from 0 to 1 for labels generated by machines, and is either 0 or 1 for labels negatively or positively verified by humans respectively. See section 3.2 for more information. We use this property to exclude images from Wake Vision if they contain only low confidence person labels.
Quality Training Set From Bounding Box Labels
@@
We allow users to adapt the Wake Vision dataset to meet the needs of specific usecases by adding several configurable options in the open-source dataset creation script. For example a configuration option exists to change whether depictions are labelled as non-persons or excluded from the dataset. A comprehensive list of Open Images labels related to Wake Vision can be found in Sec. L.
@@
Training Set Comparison
To compare the performance of the two Wake Vision training sets, we train a MobileNetV2-0.25 [26] model on each dataset for an equal number of steps and compare its performance on the common test set (Table 2). The filtered training set generated from the bounding box labels outperforms the larger training set generated from the image-level labels by 4.09% test accuracy. This indicates that label quality is more important than quantity in this setting. The gap between the two is reduced to just 0.17% when training exclusively on soft labels from a teacher model (MobileNetV2-1.0) trained on bounding box labels.
Combining the two training sets to train a model where Wake Vision (Large) acts as a pre-training set and Wake Vision (Quality) acts as a fine-tuning set achieves our best performance at a mean top-level test accuracy of 85.72%. As the images in Wake Vision (Quality) is a subset of the images in Wake Vision (Large), we do not train on more images than for Wake Vision (Large).
Table 2: Accuracy on test set trained on image-level and bounding box labels, with and without distillation. A MobileNetV2-1.0 model trained on the bounding box set is used as the teacher in KD.
Training set Label Source Training Set Size Accuracy Distilled Accuracy
Wake Vision (Large) Image-Level 5,760,428 80.80$\pm$0.1885.00$\pm$0.20%
Wake Vision (Quality) Bounding Box 1,248,230 84.89$\pm$0.1185.17$\pm$0.18%
Wake Vision (Combined) Both 5,760,428 85.72$\pm$0.04-
As our distillation results indicate, further data-centric filtering techniques may be needed to fully leverage the image-level training set. We experimented with tripling the model size and doubling the training time, but the difference in accuracy remained relatively consistent.
@@
3.2 Label Correction
Label errors are a challenge in computer vision [24, 4] that limit progress and obscures true performance. This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs. Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of its labels, aligning with best practices in the field of computer vision.
@@ 3.2 Label Correction
Given Wake Vision’s size, manually correcting label errors across the entire dataset becomes an arduous undertaking. Therefore, we prioritize labeling the Wake Vision validation and test sets so that the dataset accurately represents the model’s performance. To do so we make use of the Scale AI platform to crowd-source manual label corrections (see Appendix). With the ground truth established in the validation and test sets, Wake Vision becomes an ideal foundation for research on automated data cleaning techniques [23]. We perform an initial exploration of these methods in appendix N. We report our estimated label error rate after label corrections versus VWW in Table 3. The Wake Vision validation and test set is considerably lower than that of the VWW dataset top-level error rate.
Table 3: Estimated label error rate of VWW and Wake Vision, based on a 500 sample subset. The Wake Vision estimated error rate is broken down by set. The Wake Vision Train (Quality) and Train (Large) have different error rates due to the different sources of their labels. The Wake Vision Validation and Test sets have a lower error rate due to manual relabeling.
Dataset Wake Vision Set Label Error Rate
Visual Wake Words Train, Val, & Test 7.8%
Wake Vision Train (Large) 15.2%
Wake Vision Train (Quality) 6.8%
Wake Vision Val & Test 2.2%